Fore Farms reported a pretax operating loss of $128 million for financial reporting purposes in 2021. Contributing to the loss were (a) a penalty of $4 million assessed by the Environmental Protection Agency for violation of a federal law and paid in 2021 and (b) an estimated loss of $10 million from accruing a loss contingency. The loss will be tax deductible when paid in 2022.

The enacted tax rate is 25%. There were no temporary differences at the beginning of the year and none originating in 2021 other than those described above. Taxable income in Fores’s two previous years of operation was as follows:

2019 $ 64 million
2020 $ 40 million


Required:
1. Prepare the journal entry to recognize the income tax benefit of the net operating loss in 2021. Assume Fore will carry back its NOL to prior years.
2. What is the net operating loss reported in 2021 income statement?
3. Prepare the journal entry to record income taxes in 2022 assuming pretax accounting income is $176 million. No additional temporary differences originate in 2022.
Answer #1

Answer 1

Journal Entry
Particulars Debit Credit
Receivable income tax refund 26
Deferred tax assets 5
Income tax benifit 31
(to recognize the income tax benifit of the net operating loss in 2021)

Working note:

Receivable income tax refund = (64 + 40) * 25% = 26

Deferred tax assets = (128 - 64 - 40 - 4) * 25% = 5

Income tax benifit = ( 128 - 4) * 25% = 31

Answer 2

Income Statement 2021
Operating loss before income taxes - 128
Income tax benefit:
Benefit due to loss carryback 26
Benefit due to loss carry forward 5 31
Net income / loss - 97

Answer 3

Journal Entry
Particulars Debit Credit
Income tax expense 44
Deferred tax assets 5
Income tax payable 49
(to record income taxes in 2022)

Working note:

Income tax expense = 176 * 25% = 44

Deferred tax assets = (128 - 64 - 40 - 4) * 25% = 5

Dear Student please upvote my answer :)

Thankyou